Solve the following differential equation d^2y/dx^2 - 2dy/dx + 10y = 0,. Given Y(0) = 4. dy/dx (0) = 1
1
Expert's answer
2020-12-08T10:16:32-0500

"\frac{d^2y}{dx^2}-2\frac{dy}{dx}+10y=0\\
\text{The auxilliary equation is going to be}\\
m^2-2m+10=0\\
\text{Using quadratic formula}\\
m=\frac{-(-2)\pm\sqrt{(-2)^2-4(1)(10)}}{2(10)}\\
m=\frac{2\pm\sqrt{-36}}{2}\\
\text{m1 = 1+3i and m2 = 1- 3i}\\
Y=A_1e^{(1+3i)x}+A_2e^{(1-3i)x}\\
Y=A_1e^x.e^{3ix}+A_2e^x.e^{-3ix}\\
Y=A_1e^x[cos(3x)+isin(3x)]+A_2e^x[cos(3x)-isin(3x)]\\
Y=[(A_1+A_2)e^xcos(3x)]+[(A_1-A_2)e^xisin(3x)]\\
Let (A_1+A_2)=A\\
 (A_1-A_2)=B
\implies\\
Y=e^x[Acos(3x)+Bisin(3x)]\\
\text{since Y(0)=4, then}\\
Y(0)=e^0[Acos3(0)+Bisin3(0)]\\
\text{4=A that is}\\
\text{A = 4}\\
\frac{dy}{dx}=Ae^x[cos(3x)-3isin(3x)]+Be^x[3cos(3x)+isin(3x)]\\

\frac{dy}{dx}(0)=Ae^0[cos(3(0))-3isin(3(0))]+Be^0[3cos(3(0))+isin(3(0))]\\
\frac{dy}{dx}(0)=A+3B\\
1= 4 +3B\\
3B=-3\\
B=-1\\

Y=e^x[4cos(3x)-isin(3x)]"